# Stability analysis of action potential generation using Markov models of voltage‑gated sodium channel isoforms

### v1

@@ 1 Introduction
The distribution of sodium channel isoforms is tightly regulated across tissues, contributing to both physiological specialization and disease susceptibility. In the CNS, NaV1.1, NaV1.2, NaV1.3 and NaV1.6 predominate [24, 23]. Meanwhile, NaV1.7, NaV1.8, and NaV1.9 are expressed in the PNS [24]. NaV1.4 and NaV1.5 are primarily expressed in muscle and cardiac tissues [24, 6] NaV1.1 and NaV1.2 play an important role in excitatory and inhibitory neurons in different regions of the brain [6]. By contrast, NaV1.4 is the principal isoform in skeletal muscle, where it drives action potentials that control muscle contraction, and mutations in NaV1.4 give rise to a range of hereditary myopathies. NaV1.5, the primary cardiac isoform, is expressed in atrial, Purkinje, and ventricular cells, where it mediates the rapid depolarization of the cardiac action potential; pathogenic variants in its gene are strongly associated with life-threatening arrhythmia.
Collectively, this isoform-specific distribution highlights the complexity of sodium channel physiology and highlights the limitations of classical HH-type formulations that treat sodium conductivity as homogeneous. Accurately modeling these isoforms is essential not only for advancing our understanding of excitability and signal propagation in different tissues but also for the rational design of therapeutic interventions and for developing synthetic excitable cells. In synthetic neuroscience and bioengineering, incorporating isoform-specific dynamics will be critical for creating biologically realistic neural models and for developing targeted therapies for channelopathies and excitability disorders [2].
@@ 2 Mathematical Model
(4) $r_{ij}(V)=B^{\mathrm{hyp}}_{ij}\left[1+e^{\frac{(V-V^{\mathrm{hyp}}_{ij})}{K^{\mathrm{hyp}}_{ij}}}\Big]^{-1}+B^{\mathrm{dep}}_{ij}\Big[1+e^{\frac{(V-V^{\mathrm{dep}}_{ij})}{K^{\mathrm{dep}}_{ij}}}\right]^{-1}$
@@ 3 Methods
To explore the effect of biophysical parameters on neuronal dynamics, we perform a systematic sweep over the maximal sodium and potassium conductances $\bar{g}_{\mathrm{Na}}$ and $\bar{g}_{\mathrm{K}}$. These parameters are varies from 0 to 10 times of their physiological values. At each grid point in the parameter space, we compute the Jacobian and assess stability via its eigenvalues. We increase the stimulation current $I_{ext}$ at each point in the parameter space from 0 mA/cm2 till $50$ mA/cm2 by increments of 1 mA/cm2. We then record the minimal stimulation current needed to obtain an oscillatory solution.
@@ 4 Results
Applying the bifurcation curve identification method on the parameter space of the NaV1.1 sodium channel isoform heatmap produces figure 5. This figure shows the bifurcation structure of the Hodgkin-Huxley model in the two-dimensional parameter space defined by the maximal sodium ($\bar{g}_{\mathrm{Na}}$) and potassium ($\bar{g}_{\mathrm{K}}$) conductances. The heatmap indicates the minimal external current $I_{\mathrm{ext}}$ required to sustain a stable limit cycle at each point in the conductance space. Superimposed in red is the Hopf bifurcation curve, demarcating the boundary between quiescent and oscillatory regimes.

### v2

@@ 1 Introduction
The expression of voltage‑gated sodium channel isoforms is tightly regulated across tissues, supporting both physiological specialization and disease susceptibility. In the CNS, NaV1.1, NaV1.2, NaV1.3 and NaV1.6 are the predominant isoforms [23, 24]. By contrast, NaV1.7, NaV1.8 and NaV1.9 are largely restricted to the PNS [24]. Within the CNS, NaV1.1 and NaV1.2 play complementary roles in inhibitory and excitatory neurons, respectively [6]. In excitable non‑neuronal tissues, NaV1.4 is the principal skeletal‑muscle isoform: it drives the action potentials that trigger muscle contraction, and mutations in NaV1.4 give rise to hereditary myopathies. NaV1.5 is the primary cardiac isoform, expressed in atrial, Purkinje and ventricular cells, where it mediates the rapid depolarization of the cardiac action potential; pathogenic variants in its gene are strongly associated with life‑threatening arrhythmia [6].
Collectively, this isoform-specific distribution highlights the complexity of sodium channel physiology and highlights the limitations of classical HH-type formulations that treat sodium channel dynamics as homogeneous. Accurately modeling these isoforms is essential not only for advancing our understanding of excitability and signal propagation in different tissues but also for the rational design of therapeutic interventions and for developing synthetic excitable cells. In synthetic neuroscience and bioengineering, incorporating isoform-specific dynamics will be critical for creating biologically realistic neural models and for developing targeted therapies for channelopathies and excitability disorders [2].
@@ 2 Mathematical Model
(4) $r_{ij}(V)=B^{\mathrm{hyp}}_{ij}\left[1+e^{\frac{(V-V^{\mathrm{hyp}}_{ij})}{K^{\mathrm{hyp}}_{ij}}}\right]^{-1}+B^{\mathrm{dep}}_{ij}\left[1+e^{\frac{(V-V^{\mathrm{dep}}_{ij})}{K^{\mathrm{dep}}_{ij}}}\right]^{-1}$
@@ 3 Methods
To explore the effect of biophysical parameters on neuronal dynamics, we perform a systematic sweep over the maximal sodium and potassium conductances $\bar{g}_{\mathrm{Na}}$ and $\bar{g}_{\mathrm{K}}$. These parameters are varied from 0 to 10 times of their physiological values. At each grid point in the parameter space, we compute the Jacobian and assess stability via its eigenvalues. We increase the stimulation current $I_{ext}$ at each point in the parameter space from 0 mA/cm2 till $50$ mA/cm2 by increments of 1 mA/cm2. We then record the minimal stimulation current needed to obtain an oscillatory solution.
@@ 4 Results
Applying bifurcation analysis to the hybrid model comprising the NaV1.1 Sodium Markov Model scheme and the KV3.1 Hodgkin–Huxley potassium channel yields the heatmap in Fig. 5. This figure illustrates the dynamical structure of the system across the two‑dimensional parameter space spanned by the maximal sodium ($\bar{g}_{\mathrm{Na}}$) and potassium ($\bar{g}_{\mathrm{K}}$) conductances. Each point is colored by the minimal external current $I_{\mathrm{ext}}$ needed to sustain a stable limit cycle, with darker shades indicating higher thresholds or absence of oscillations. A red curve superimposed on the heatmap marks the Hopf bifurcation boundary separating spontaneous and stimulated oscillation regimes.